OpenStudy (anonymous):

find the value of k so that the remainder of (-kx^4 - 146x^2 + 32)divided by (x-4)is 0

OpenStudy (anonymous):

(-kx^4 - 146x^2 + 32)/(x-4)=0 is u meaning?

OpenStudy (anonymous):

yes but u solve for k

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let t=x^2 then -kt^2-146t+32=0and t\[\ge0\] \[\Delta \ge0 and t \ge0 \]then see t=o is not solution, so k=(32-146t)/t^2 dk/dt=(146t^2-64t)t^4 =0 =>t=0 or t=64/146 at t=32/73 k min =>k\[\ge k(32/73)=\]